OpenStudy (anonymous):

The position function for an object is s(t)=3t^2+2t+5 where s is measured in meters and t is measured in seconds. Find the velocity of the object when t=2. If the radius of a right circular cone equals its height, how fast is the volume changing with respect to changes in r when r=2 feet?

OpenStudy (mathmate):

Given \(s(t)=3t^2+2t+5\) then \(s(2)=3(2)^2+2(2)+5 = ?\) Given r=h, \(\large V=\frac{\pi r^3}{3}\) \(\large \frac{dV}{dr}=V'(r)=\pi r^2\) Calculate \(V'(2)\) for rate of change of volume at r=2.

OpenStudy (anonymous):

thank you

OpenStudy (mathmate):

You're welcome! :)